OpenStudy (anonymous):

Will fan and medal :) please help!! Choose all answers that are correct. 1. (1 pt) Which literal equations are equivalent to p = mv? Choose all answers that are correct. A. m=p/v B.m= v/p C. v= m/p D. v=p/m

OpenStudy (anonymous):

\[\huge \color{red}p=\color{blue}m\color{green}v\] \[\huge \color{red}{12}=\color{blue}4\times \color{green}3\]

OpenStudy (anonymous):

\[\huge \frac{\color{red}{12}}{\color{blue}4}=\color{green}3\]

OpenStudy (anonymous):

D?

OpenStudy (anonymous):

\[\huge \frac{\color{red}{12}}{\color{green}3}=\color{blue} 4\]

OpenStudy (anonymous):

A &D

OpenStudy (anonymous):

yeah D for sure

OpenStudy (anonymous):

and A right

OpenStudy (anonymous):

Thank you for your help :)

OpenStudy (anonymous):

yw